OpenStudy (kyle755):

Solve for x: 2 over 5 (x − 2) = 4x. A. 2/9 B. 9 c. negative 2/9 D. negative 9/2

OpenStudy (kyle755):

The second one

OpenStudy (confluxepic):

Is that \(\frac{ 2 }{ 5(x-2)}\) or \(\frac{ 2 }{ 5 }(x-2)\)?

OpenStudy (confluxepic):

\(\Huge\frac{ 2 }{ 5 }(x-2)=4x\)

OpenStudy (confluxepic):

Just manipulate the equations. Simple. :)

OpenStudy (confluxepic):

First simplify them.

OpenStudy (confluxepic):

Get rid of the fraction to make life easier. Do you know how to do that? @Kyle755

OpenStudy (kyle755):

yes 2/5 * 5/2

OpenStudy (confluxepic):

Good thinking. I like you. Alright, what do you get after you multiply it by 5/2. (Remember to multiply both sides by the same thing.)

OpenStudy (kyle755):

1

OpenStudy (confluxepic):

\(\ (x-2)=4x \times (5/2)\) Is that correct?

OpenStudy (kyle755):

no

OpenStudy (confluxepic):

Why not? I multiplied both sides of the equation by 5/2.

OpenStudy (kyle755):

it would be yes

OpenStudy (confluxepic):

\(\ x-2=4x \times (5/2)\) What is 4x times 5/2?

OpenStudy (kyle755):

6

OpenStudy (confluxepic):

\(\Huge\ I~am~going~to~solve~only~the\\\Huge\ right~side~of~the~equation.\) \(\Huge\ 4x \times \frac{5}{2}=\) \(\Huge\ \frac{4x\times 5}{2}=\) \(\Huge\ \frac{20x}{2}=\) \(\Huge\ 10x\) \(\Huge\ x-2=10x\)

OpenStudy (confluxepic):

Can you do the rest or do you need more of my help?

OpenStudy (kyle755):

ok and thank you and I can do the rest

OpenStudy (confluxepic):

You're welcome. :)